### v1

@@ Abstract
Model predictive control (MPC) is widely used in process control due to its interpretability and ability to handle constraints. As a parametric policy in reinforcement learning (RL), MPC offers strong initial performance and low data requirements compared to black-box policies like neural networks. However, most RL methods rely on first-order updates, which scale well to large parameter spaces but converge at most linearly, making them inefficient when each policy update requires solving an optimal control problem, as is the case with MPC. While MPC policies are typically sparsely parameterized and thus amenable to second-order approaches, existing second-order methods demand second-order policy derivatives, which can be computationally and memory-wise intractable.
This work introduces a Gauss-Newton approximation of the deterministic policy Hessian that eliminates the need for second-order policy derivatives, enabling superlinear convergence with minimal computational overhead. To further improve robustness, we propose a momentum-based Hessian averaging scheme for stable training under noisy estimates. We demonstrate the effectiveness of the approach on a nonlinear continuously stirred tank reactor (CSTR), showing faster convergence and improved data efficiency over state-of-the-art first-order methods.
@@ 1 Introduction
Model predictive control (MPC) is a popular approach for the control of chemical processes because it can deal with highly dynamic, nonlinear multivariable systems subject to safety and quality constraints. In MPC, a system model of the controlled plant is used to predict and optimize the future behavior of the system states so that a objective function is minimized while ensuring process constraints. The first control action of the optimal solution is then deployed to the real plant, the next state is measured or estimated, and the process is repeated. Since MPC delivers a sequence of optimal control inputs and a sequence of upcoming states, the control policy is highly interpretable. As an established method, many theoretical results regarding stability and feasibility exist (Rawlings et al., 2024).
@@
Besides rigorous model-based optimal control approaches like MPC, there are emerging data-driven approaches such as reinforcement learning (RL) that show suitability and success in the process industry (Yoo et al., 2021; Gopaluni et al., 2020; Lee and Lee, 2007). Model-free RL as a subfield RL focuses on learning an optimal policy through trial-and-error interaction between an agent (controller) and its environment (plant), without relying on a model of the environment. In practice, this optimal policy is often approximated directly using a rich parametric function approximator such as neural networks (NN). Contrary to MPC, this has the benefit that no differentiable system model is necessary to perform online optimization, which is especially interesting if models are hard to build or impossible to differentiate. However, in general model-free RL approaches need large amounts of data, which can be prohibitive for application to industrial scale processes (Recht, 2019).
@@
We focus in this work on two main reasons for the large data requirements of RL: the difficulty of initialization and the use of first-order optimization approaches. If expert knowledge is available beforehand, the NN policy of the RL agent can be pretrained to delivers a reasonable initial performance. However, if expert knowledge is not available, the NN policy is often initialized with random values such that many updates are necessary to achieve a similar performance. Most popular model-free RL algorithms only consider update rules that consist of first-order derivative information such as Fujimoto et al. (2018) and Haarnoja et al. (2018). In either case, the maximum convergence rate of first-order approaches is limited to be linear (Bottou et al., 2018). Methods that consider (approximate) second-order derivative information such as Newton, quasi-Newton, or inexact Newton methods can provide up to quadratic convergence. However, all those approaches must store at least a Hessian matrix approximation in memory and a linear system of equations must be solved for each update step at least approximately. This can be prohibitive especially when rich parameterized policy approximators such as NNs are used. Works such as Korbit et al. (2024), Liu et al. (2024) and Yao et al. (2021) address the memory issue by approximating the Hessian via scalars or diagonal representations, whereas works such as Chauhan et al. (2020) and Schraudolph et al. (2007) only solve the linear system of equations inexactly.
@@
Recent research identifies the challenges described above and proposes to use MPC instead of NNs as a policy for the RL agent (Gros and Zanon, 2020; Banker and Mesbah, 2025; Brandner and Lucia, 2024; Mayfrank et al., 2024). Figure 1 illustrates how NN policies can be replaced by MPC. Reiter et al. (2025) provide a comprehensive review for MPC used as policy in RL. The fundamental idea is to introduce tunable parameters to the MPC’s optimal control problem and to adjust them in the same way as the parameters in NN policies. Contrary to NN policies, MPC provides a reasonable initial policy since the proposed control action results from the solution of a model-based optimal control approach. Further, MPC typically has fewer tuning parameters than NN policies. Due to the sparse parameterization of MPC policies, Hessian computation becomes tractable, allowing more sophisticated updates with superlinear convergence rates, which are of significant interest for MPC policies because each RL transition requires the solution of an optimization problem, which is computationally expensive.
@@
Different update rules were proposed in the past to improve the data efficiency. Kakade (2001) introduces the natural policy gradient, which scales the policy gradient by the inverse of the Fisher information matrix. Although being limited to linear convergence, it often shows overall practical improvement. Further improvement is obtained via quasi-Newton approaches with a superlinear convergence rate as proposed in Furmston et al. (2016), Jha et al. (2020), and Kordabad et al. (2022). With a special focus on MPC policies, Brandner and Lucia (2024) propose a tailored algorithm that reduces the number of updates significantly. However, the second-order policy derivative with respect to its parameters needs to be computed. Brandner and Lucia (2024) show that the second-order derivative of MPC policies can be obtained by the solution of a linear system of equations. Unfortunately, the right-hand side of the linear system involves third-order derivatives of the MPC’s Lagrangian with respect to all primal-dual variables, which can be computationally demanding and memory intensive, especially for high dimensional and highly nonlinear system models.
Lastly, it needs to be said that all considered update rules, regardless if first or second-order, require to compute expected values over the distribution of states that appear under closed-loop operation, and further assume perfect knowledge of the action-value function, which evaluates how good a certain action in a given state is, when following the present policy. In practice however, these expected values can typically only be estimated empirically by a sufficiently large set of closed-loop samples, whereas the action-value function can only be approximated. Both factors contribute to the fact that the gradient and the Hessian cannot be computed perfectly, which can lead to unstable training especially if the Hessian eigenvalues are estimated wrongly.
In this work, we address both the high memory requirements and high computational demand when using MPC policies in second-order RL due to the second-order policy derivatives, and the training instabilities due to poor gradient and Hessian estimation. Our main contributions are twofold:
@@
    We propose a Gauss-Newton approximation of the deterministic policy Hessian that does not require second-order policy derivatives, leading to significant computational savings. We rigorously prove that our approach can still lead to superlinear convergence despite the simpler required computations. Our proposed approximation alleviates the high data requirements compared to the established first-order policy updates in RL.
@@
    We derive a momentum-based update rule for Hessian estimation that returns an unbiased exponentially averaged Hessian estimate. This estimate is robust with respect to noise in local Hessian samples and therefore enables stable and fast training as it reduces the noise in the eigenvalues of the Hessian. A reliable Hessian estimate is very relevant for MPC policies in RL because the influence of the policy parameters on the objective can typically vary by several orders of magnitude.
The work is structured as follows. We introduce the relevant preliminaries on Markov decision processes, iterative policy optimization via RL, and MPC as policy approximator in RL in Section 2. We present the proposed Gauss-Newton Hessian approximation and momentum-based update rule in Section 3. We illustrate the superlinear convergence capability of the Gauss-Newton approximation on an analytical case study in Section 4 and show the practical improvement of the Gauss-Newton approximation and the proposed momentum-based update rule at the example of a nonlinear continuously stirred tank reactor (CSTR) in Section 5. Section 6 concludes the work and provides perspectives for future work.
@@ 2.1 Markov decision processes
RL can solve problems that can be formalized as a Markov decision process (MDP). MDPs are a general mathematical modeling framework for sequential decision making (Sutton and Barto, 2018). In an MDP, an agent (controller) interacts with an environment (plant) as depicted in Figure 1. The agent computes a certain action $\boldsymbol{a}\in\mathcal{A}\subseteq\mathbb{R}^{n_{\boldsymbol{a}}}$ according to the agent’s decision making policy $\pi$, given the current environmental state $\boldsymbol{s}\in\mathcal{S}\subseteq\mathbb{R}^{n_{\boldsymbol{s}}}$. In this work, we focus on deterministic polices $\pi:\mathcal{S}\rightarrow\mathcal{A}$ that directly map the current state $\boldsymbol{s}$ to its proposed action $\boldsymbol{a}$. This action $\boldsymbol{a}$ is handed to the environment, which transitions from the current state $\boldsymbol{s}$ to a subsequent state $\boldsymbol{s}^{+}\in\mathcal{S}$ according to the MDP’s conditional transition probability $\rho:\mathcal{S}\times\mathcal{S}\times\mathcal{A}\rightarrow[0,\infty)$. The subsequent state is sampled from $\rho$ according to
@@
Once that the environment transitioned from state $\boldsymbol{s}$ to $\boldsymbol{s}^{+}$, the new environmental state is handed to the agent together with a reward $r\in\mathbb{R}$. This reward $r$ is computed according the environment’s reward function $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, which depends on the current state $\boldsymbol{s}$ and the taken action $\boldsymbol{a}$, and serves as an immediate measure of how good the current state-action pair $\boldsymbol{s},\boldsymbol{a}\right>$ is, while neglecting future consequences due to upcoming subsequent states.
@@
[…]
@@ 2.2 Iterative policy optimization in reinforcement learning
To find the policy $\pi^{\star}$ that solves (6), an infinite dimensional optimization problem over the space of all admissible policy functions must be solved. This is intractable in most cases. Therefore, the policy is often parameterized by a function approximator $\pi_{\boldsymbol{\theta}}(\boldsymbol{s})$, where $\boldsymbol{\theta}\in\mathbb{R}^{n_{\boldsymbol{\theta}}}$ are tunable policy parameters (Sutton and Barto, 2018). The expected cumulative reward transforms from a function of the policy $J(\pi)$ into a function of the policy parameters $J(\pi_{\boldsymbol{\theta}})=J(\boldsymbol{\theta})$. The infinite dimensional optimization problem (6) simplifies to the finite dimensional optimization problem
@@
In its general form, optimization problem (9) can be categorized as an unconstrained nonlinear program with stochastic objective since $J$ is defined in terms of expected values as shown in (5). Iterative optimization algorithms are standard tools to solve such optimization problems numerically (Nocedal and Wright, 2006). The basic idea is to start at an initial guess $\boldsymbol{\theta}_{0}$ of the optimal solution and improve the guess iteratively. The general iterative update process at the $k$-th iteration step can be described as
@@
[…]
@@ 2.2.1 First-order policy optimization
An optimization algorithm is considered to be of first-order, if it only uses first-order derivative information of the optimization problem during the optimization process. In case of (9), this refers to the gradient of the expected cumulative reward with respect to the policy parameters $\nabla_{\boldsymbol{\theta}}J(\boldsymbol{\theta})\in\mathbb{R}^{n_{\boldsymbol{\theta}}}$. In order to preserve a concise and clutter-free notation, we introduce the gradient of the objective function, evaluated at the current iterate $\boldsymbol{\theta}_{k}$ as $\boldsymbol{g}_{k}=\nabla_{\boldsymbol{\theta}}J(\boldsymbol{\theta})|_{\boldsymbol{\theta}=\boldsymbol{\theta}_{k}}$. Silver et al. (2014) derive the deterministic policy gradient as
@@
where $\nabla_{\boldsymbol{\theta}}\pi_{\boldsymbol{\theta}}(\boldsymbol{s})\in\mathbb{R}^{n_{\boldsymbol{a}}\times n_{\boldsymbol{\theta}}}$ denotes the Jacobian of the policy with respect to its parameters and $\nabla_{\boldsymbol{a}}Q^{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s},\boldsymbol{a})\in\mathbb{R}^{n_{\boldsymbol{a}}}$ denotes the gradient of the policy’s action-value function with respect to the taken actions, which is evaluated at the policy’s proposed action $\boldsymbol{a}=\pi_{\boldsymbol{\theta}}(\boldsymbol{s})$. The expected value is taken over all $\boldsymbol{s}$ that come up during closed-loop operation.
As the simplest form of a first-order method for stochastic objectives, we introduce the stochastic gradient ascent algorithm (Robbins and Monro, 1951). The stochastic gradient ascent algorithm stems from the deterministic gradient ascent algorithm, which updates the current iterate $\boldsymbol{\theta}_{k}$ into the direction of steepest ascent, which is the gradient of the objective function at the current iterate $\boldsymbol{g}_{k}$. Since the deterministic policy gradient (12) is computed using an expected value, it can only be estimated empirically using a finite amount of samples $N\in\mathbb{N}$ such that
$\tilde{\boldsymbol{g}}_{k}=\frac{1}{N}\sum_{i=1}^{N}\nabla_{\boldsymbol{\theta}}\pi_{\boldsymbol{\theta}}^{\top}(\boldsymbol{s}_{i})\nabla_{\boldsymbol{a}}Q^{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s}_{i},\boldsymbol{a})|_{\boldsymbol{a}=\pi_{\boldsymbol{\theta}}(\boldsymbol{s}_{i})}$ (13)
@@
Many works aim to improve the stability and the convergence of stochastic gradient ascent by applying an exponential moving average to observed gradient estimates (Kingma and Ba, 2015; Duchi et al., 2011; Rumelhart et al., 1986). These methods are called momentum-based methods. In its core, momentum-based methods update an exponential moving average $\boldsymbol{m}_{k}\in\mathbb{R}^{n_{\boldsymbol{\theta}}}$ of past gradients with the newly obtained gradient estimate $\tilde{\boldsymbol{g}}_{k}$ according to
$\boldsymbol{m}_{k}=\beta\boldsymbol{m}_{k-1}+(1-\beta)\tilde{\boldsymbol{g}}_{k}$ (15)
where $0\leq\beta\leq 1$ is an exponential decay factor. The larger $\beta$ the slower the moving average $\boldsymbol{m}_{k}$ adapts to the current gradient estimate. Typically, the moving average is initialized as a zero vector $\boldsymbol{m}_{0}=0$. This causes initialization bias of the exponential moving average towards zero, which results in slow improvement in early stages of the optimization. Kingma and Ba (2015) correct the initialization bias by
$\hat{\boldsymbol{m}}_{k}=\frac{\boldsymbol{m}_{k}}{(1-\beta^{k})}$ (16)

### v2

@@ Abstract
Model predictive control (MPC) is widely used in process control due to its interpretability and ability to handle constraints. As a parametric policy in reinforcement learning (RL), MPC offers strong initial performance and low data requirements compared to black-box policies like neural networks. However, most RL methods rely on first-order updates, which scale well to large parameter spaces but converge at most linearly, making them inefficient when each policy update requires solving an optimal control problem, as is the case with MPC. While MPC policies are typically low parameterized and thus amenable to second-order approaches, existing second-order methods demand second-order policy derivatives, which can be computationally intractable.
This work introduces a Gauss-Newton approximation of the deterministic policy Hessian that eliminates the need for second-order policy derivatives, enabling superlinear convergence with minimal computational overhead. To further improve robustness, we propose a momentum-based Hessian averaging scheme for stable training under noisy estimates coupled with an adaptive trust-region. We demonstrate the effectiveness of the approach on a nonlinear continuously stirred tank reactor (CSTR), showing faster convergence and improved data efficiency over state-of-the-art first-order methods and deep RL approaches.
@@ 1 Introduction
Model predictive control (MPC) is a popular approach for the control of highly dynamic, nonlinear multivariable systems subject to safety and quality constraints. In MPC, a control plant system model is used to predict and optimize the future behavior of the system states so that an objective function is minimized while enforce process constraints. The first control action of the optimal solution is then deployed to the real plant, the next state is measured or estimated, and the process is repeated. Since MPC delivers a sequence of optimal control inputs and a sequence of upcoming states, the control policy is highly interpretable. As an established method, many theoretical results regarding stability and feasibility exist (rawlingsModelPredictiveControl2017).
@@
Besides rigorous model-based optimal control approaches like MPC, there are emerging data-driven approaches such as reinforcement learning (RL) that show suitability and success in the process industry (yooReinforcementLearningBatch2021; gopaluniModernMachineLearning2020; leeIterativeLearningControl2007). Model-free RL as a subfield RL focuses on learning an optimal policy through trial-and-error interaction between an agent (controller) and its environment (plant), without relying on a model of the environment. In practice, this optimal policy is often approximated directly using a rich parametric function approximator such as neural networks (NN). Contrary to MPC, this has the benefit that no differentiable system model is necessary to perform online optimization, which is especially interesting if models are hard to build or impossible to differentiate. However, in general model-free RL approaches need large amounts of data, which can be prohibitive for application to industrial scale processes (rechtTourReinforcementLearning2019).
@@
We focus in this work on two main reasons for the large data requirements of RL: the difficulty of initialization and the use of first-order optimization approaches. If expert knowledge is available beforehand, the NN policy of the RL agent can be pretrained to deliver a reasonable initial performance. However, if expert knowledge is not available, the NN policy is often initialized with random values such that many updates are necessary to reach an acceptable performance. Most popular model-free RL algorithms only consider update rules that consist of first-order derivative information such as fujimotoAddressingFunctionApproximation2018 and haarnojaSoftActorCriticOffPolicy2018. In either case, the maximum convergence rate of first-order approaches is limited to be linear (bottouOptimizationMethodsLargeScale2018). Methods that consider (approximate) second-order derivative information such as Newton, quasi-Newton, or inexact Newton methods can provide up to quadratic convergence. However, all those approaches must store at least a Hessian matrix approximation in memory and a linear system of equations must be solved for each update step at least approximately. This can be prohibitive especially when rich parameterized policy approximators such as NNs are used. Works such as korbitExactGaussNewtonOptimization2024, liuSophiaScalableStochastic2024 and yaoADAHESSIANAdaptiveSecond2021 address the memory issue by approximating the Hessian via scalars or diagonal representations, whereas works such as chauhanStochasticTrustRegion2020 and schraudolphStochasticQuasinewtonMethod2007 only solve the linear system of equations inexactly.
@@
Recent research identifies the challenges described above and proposes to use MPC instead of NNs as a policy for the RL agent (grosDataDrivenEconomicNMPC2020; bankerGradientBasedFrameworkBilevel2025; brandnerReinforcedModelPredictive2024b; mayfrankEndtoendReinforcementLearning2024). Figure 1 illustrates how NN policies can be replaced by MPC. reiterSynthesisModelPredictive2025 provide a comprehensive review for MPC used as policy in RL. The fundamental idea is to introduce tunable parameters to the MPC’s optimal control problem and to adjust them in the same way as the parameters in NN policies. Contrary to NN policies, MPC provides a reasonable initial policy by solving a model-based optimal control problem that directly converts system knowledge into control actions. It also provides a universal approximator to the solution of Markov decision processes (MDP), which states the theoretical framework for RL (grosDataDrivenEconomicNMPC2020). Further, MPC typically has fewer tuning parameters than NN policies. Due to the low parameterization of MPC policies, Hessian computation becomes tractable, allowing more sophisticated updates with superlinear convergence rates, which are of significant interest for MPC policies because each RL transition requires the solution of an optimization problem, which is computationally expensive.
@@
Different update rules were proposed in the past to improve data efficiency. Kakade_NaturalPolicyGradient_2001 introduces the natural policy gradient, which scales the policy gradient by the inverse of the Fisher information matrix. Although being limited to linear convergence, it often shows overall practical improvement. Further improvement is obtained via quasi-Newton approaches with a superlinear convergence rate as proposed in furmstonApproximateNewtonMethods2016, jhaQuasiNewtonTrustRegion2020, and kordabadQuasiNewtonIterationDeterministic2022. With a special focus on MPC policies, brandnerReinforcedModelPredictive2024b propose a tailored algorithm that reduces the number of updates significantly. However, the second-order policy derivative of the MPC policy with respect to its parameters needs to be computed. brandnerReinforcedModelPredictive2024b show that the second-order derivative of MPC policies can be obtained by the solution of a linear system of equations. Unfortunately, the right-hand side of the linear system involves third-order derivatives of the MPC’s Lagrangian with respect to all primal-dual variables, which can be computationally demanding and memory intensive, especially for high dimensional and highly nonlinear system models.
Lastly, all considered update rules, regardless if first or second-order, require to compute expected values and further assume perfect knowledge of the action-value function, which evaluates how good a certain action in a given state is, when following the present policy. In practice however, these expected values and the action-value function can only be approximated. Both factors contribute to the fact that the gradient and the Hessian cannot be computed perfectly, which can lead to unstable training.
In this work, we address both the high memory requirements as well as the high computational demand when using MPC policies in second-order RL due to the second-order policy derivatives, and the training instabilities due to poor gradient and Hessian estimation. Our main contributions are twofold:
@@
    We propose a Gauss-Newton approximation of the deterministic policy Hessian that does not require second-order policy derivatives, leading to significant computational savings. We rigorously prove that our approach can still lead to superlinear convergence due to faster contraction despite the simpler required computations. Our proposed approximation alleviates the high data requirements compared to the established first-order policy updates in RL.
@@
    We derive a momentum-based update rule for Hessian estimation that returns an unbiased exponentially averaged Hessian estimate. This estimate is robust with respect to noise in local Hessian samples and therefore enables stable and fast training as it reduces the noise in the eigenvalues of the Hessian. A reliable Hessian estimate is very relevant for MPC policies in RL because the influence of the policy parameters on the objective can typically vary by several orders of magnitude. We embed this unbiased exponentially averaged Hessian estimate of the Gauss-Newton approximation within an adaptive trust-region constrained optimization problem to compute robust parameter updates.
The work is structured as follows. We introduce the relevant preliminaries on Markov decision processes, iterative policy optimization via RL, and MPC as policy approximator in RL in Section 2. We present the proposed Gauss-Newton Hessian approximation, the momentum-based update rule, and the adaptive trust-region constrained update computation in Section 3. We illustrate the superlinear convergence capability of the Gauss-Newton approximation on an analytical case study in Section 4 and show the practical improvement of the proposed approach at the example of a nonlinear continuously stirred tank reactor (CSTR) in Section 5. Section 6 concludes the work and provides perspectives for future work.
@@ 2.1 Markov decision processes
RL can solve problems that can be formalized as a MDP. MDPs are a general mathematical modeling framework for sequential decision making (suttonReinforcementLearningIntroduction2018). In an MDP, an agent (controller) interacts with an environment (plant) as depicted in Figure 1. The agent computes a certain action $\boldsymbol{a}\in\mathcal{A}\subseteq\mathbb{R}^{n_{\boldsymbol{a}}}$ according to the agent’s decision making policy $\pi$, given the current environmental state $\boldsymbol{s}\in\mathcal{S}\subseteq\mathbb{R}^{n_{\boldsymbol{s}}}$. In this work, we focus on deterministic polices $\pi:\mathcal{S}\rightarrow\mathcal{A}$ that directly map the current state $\boldsymbol{s}$ to its proposed action $\boldsymbol{a}$. This action $\boldsymbol{a}$ is handed to the environment, which transitions from the current state $\boldsymbol{s}$ to a subsequent state $\boldsymbol{s}^{+}\in\mathcal{S}$ according to the MDP’s conditional transition probability $\rho:\mathcal{S}\times\mathcal{S}\times\mathcal{A}\rightarrow[0,\infty)$. The subsequent state is sampled from $\rho$ according to
@@
Once that the environment transitioned from state $\boldsymbol{s}$ to $\boldsymbol{s}^{+}$, the new environmental state is handed to the agent together with a reward $r\in\mathbb{R}$. This reward $r$ is computed according the environment’s reward function $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, which depends on the current state $\boldsymbol{s}$ and the taken action $\boldsymbol{a}$, and serves as an immediate measure of how good the current state-action pair $\left(\boldsymbol{s},\boldsymbol{a}\right)$ is, while neglecting future consequences due to upcoming subsequent states.
@@
$\begin{aligned} V^{\pi}(\boldsymbol{s})=&R(\boldsymbol{s},\pi(\boldsymbol{s}))+\gamma\mathbb{E}_{\boldsymbol{s}^{+}}\left[R(\boldsymbol{s}^{+},\pi(\boldsymbol{s}^{+}))\right]\\ &+\gamma^{2}\mathbb{E}_{\boldsymbol{s}^{+}}\left[\mathbb{E}_{\boldsymbol{s}^{++}}\left[R(\boldsymbol{s}^{++},\pi(\boldsymbol{s}^{++}))\right]\right]+\ldots\end{aligned}$
@@ 2.2 Iterative policy optimization in reinforcement learning
To find the policy $\pi^{\star}$ that solves (6), an infinite dimensional optimization problem over the space of all admissible policy functions must be solved. This is intractable in most cases. Therefore, the policy is often parameterized by a function approximator $\pi_{\boldsymbol{\theta}}(\boldsymbol{s})$, where $\boldsymbol{\theta}\in\mathbb{R}^{n_{\boldsymbol{\theta}}}$ are tunable policy parameters (suttonReinforcementLearningIntroduction2018). The expected cumulative reward transforms from a function of the policy $J(\pi)$ into a function of the policy parameters $J(\pi_{\boldsymbol{\theta}})=J(\boldsymbol{\theta})$. The infinite dimensional optimization problem (6) simplifies to the finite dimensional optimization problem
@@
In its general form, optimization problem (9) can be categorized as an unconstrained nonlinear program with stochastic objective since $J$ is defined in terms of expected values as shown in (5). Iterative optimization algorithms are standard tools to solve such optimization problems numerically (nocedalNumericalOptimization2006). The basic idea is to start at an initial guess $\boldsymbol{\theta}_{0}$ of the optimal solution and improve the guess iteratively. The general iterative update process at the $k$-th iteration step can be described as
@@
where $\|\cdot\|$ denotes a vector-norm and $\epsilon\geq 0$ is a constant.
@@ 2.2.1 First-order policy optimization
An optimization algorithm is considered to be of first-order if it only uses first-order derivative information of the optimization problem during the optimization process. In case of (9), this refers to the gradient of the expected cumulative reward with respect to the policy parameters $\nabla_{\boldsymbol{\theta}}J(\boldsymbol{\theta})\in\mathbb{R}^{n_{\boldsymbol{\theta}}}$. In order to preserve a concise and clutter-free notation, we introduce the gradient of the objective function, evaluated at the current iterate $\boldsymbol{\theta}_{k}$ as $\boldsymbol{g}_{k}=\nabla_{\boldsymbol{\theta}}J(\boldsymbol{\theta})|_{\boldsymbol{\theta}=\boldsymbol{\theta}_{k}}$. silverDeterministicPolicyGradient2014 derive the deterministic policy gradient as
@@
where $\nabla_{\boldsymbol{\theta}}\pi_{\boldsymbol{\theta}}(\boldsymbol{s})\in\mathbb{R}^{n_{\boldsymbol{a}}\times n_{\boldsymbol{\theta}}}$ denotes the Jacobian of the policy with respect to its parameters and $\nabla_{\boldsymbol{a}}Q^{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s},\boldsymbol{a})\in\mathbb{R}^{n_{\boldsymbol{a}}}$ denotes the gradient of the policy’s action-value function with respect to the taken actions, which is evaluated at the policy’s proposed action $\boldsymbol{a}=\pi_{\boldsymbol{\theta}}(\boldsymbol{s})$. The expectation $\mathbb{E}_{\boldsymbol{s}}[\cdot]$ denotes the expectation over the discounted sum of the Markov chain that occurs under closed-loop operation with policy $\pi_{\boldsymbol{\theta}}$.
As the simplest form of a first-order method for stochastic objectives, we introduce the stochastic gradient ascent algorithm (robbinsStochasticApproximationMethod1951). The stochastic gradient ascent algorithm stems from the deterministic gradient ascent algorithm, which updates the current iterate $\boldsymbol{\theta}_{k}$ into the direction of steepest ascent, which is the gradient of the objective function at the current iterate $\boldsymbol{g}_{k}$. Since the deterministic policy gradient (12) is computed using an expected value, it can only be estimated empirically using a finite amount of samples from $N_{\mathrm{IC}}\in\mathbb{N}$ initial conditions, each having episodes a length of $N_{\mathrm{ep},i}\in\mathbb{N}$ such that
$\tilde{\boldsymbol{g}}_{k}=\frac{1}{N_{\mathrm{IC}}}\sum_{i=1}^{N_{\mathrm{IC}}}\sum_{j=0}^{N_{\mathrm{ep},i}}\gamma^{j}\nabla_{\boldsymbol{\theta}}\pi_{\boldsymbol{\theta}}^{\top}(\boldsymbol{s}_{i,j})\nabla_{\boldsymbol{a}}Q^{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s}_{i,j},\boldsymbol{a})|_{\boldsymbol{a}=\pi_{\boldsymbol{\theta}}(\boldsymbol{s}_{i,j})}$ (13)
@@
Many works aim to improve the stability and the convergence of stochastic gradient ascent by applying an exponential moving average to the sampled gradient estimates (kingmaAdamMethodStochastic2015; Duchi_2011_AdaGrad; rumelhartLearningRepresentationsBackpropagating1986). These methods are called momentum-based methods. In its core, momentum-based methods update an exponential moving average $\boldsymbol{m}_{k}\in\mathbb{R}^{n_{\boldsymbol{\theta}}}$ of past gradients with the newly obtained gradient estimate $\tilde{\boldsymbol{g}}_{k}$ according to
$\boldsymbol{m}_{k}=\beta_{1}\boldsymbol{m}_{k-1}+(1-\beta_{1})\tilde{\boldsymbol{g}}_{k}$ (15)
where $0\leq\beta_{1}\leq 1$ is an exponential decay factor. The larger $\beta_{1}$, the slower the moving average $\boldsymbol{m}_{k}$ adapts to the current gradient estimate. Typically, the moving average is initialized as a zero vector $\boldsymbol{m}_{0}=0$. This causes initialization bias of the exponential moving average towards zero, which results in slow improvement in early stages of the optimization. kingmaAdamMethodStochastic2015 correct the initialization bias by
$\hat{\boldsymbol{m}}_{k}=\frac{\boldsymbol{m}_{k}}{(1-\beta_{1}^{k})}$ (16)
@@
The entries of the bias-corrected exponential moving gradient average $\hat{\boldsymbol{m}}_{k}$ can still range over several orders of magnitude, which can cause rapid learning in directions with large components and small updates into directions with small entries. The state-of-the-art first-order optimizer Adam (kingmaAdamMethodStochastic2015) tackles this challenge by rescaling all components by an exponential moving average of the magnitude of each individual component. For this, the second exponential moving average $\boldsymbol{v}_{k}\in\mathbb{R}^{n_{\boldsymbol{\theta}}}$ is introduced. It captures the square of each gradient entry as
$\boldsymbol{v}_{k}=\beta_{2}\boldsymbol{v}_{k-1}+(1-\beta_{2})\tilde{\boldsymbol{g}}_{k}\odot\tilde{\boldsymbol{g}}_{k}$ (18)
with $0\leq\beta_{2}\leq 1$ and $\odot$ denoting the Hadamard (element-wise) product. The introduced initialization bias is corrected similar to (16). Finally, the improved update rule of the Adam optimizer is defined as
$\boldsymbol{p}_{k}=\alpha\frac{\hat{\boldsymbol{m}}_{k}}{\sqrt{\hat{\boldsymbol{v}}_{k}}+\epsilon}$ (19)
where $\epsilon>0$ is a constant, introduced for numerical stability, and the division operation is meant to be element-wise.